OpenStudy (juan1857):

Differentiate the function with respect to x: y=((x+5)^5-1)^1

OpenStudy (juan1857):

\[y=((x+5)^{5}-1)^{4}\]

Nnesha (nnesha):

chain rule ?? :)

OpenStudy (juan1857):

\[4((x+5)^{5}-1)\]

OpenStudy (juan1857):

do I use the power rule next?

OpenStudy (juan1857):

wait it should be \[4(((x+5)^{5})-1)^{3}\]

OpenStudy (juan1857):

i think...

OpenStudy (mathmale):

Are you sure that your expression ends in ^1? Any base, other than 0, raised to the power 1, is equal to the base itself. Check to ensure you have copied the problem correctly.

OpenStudy (3mar):

What about if you divide it into pieces?

OpenStudy (juan1857):

i originally wrote it wrong @mathmale

Nnesha (nnesha):

yes ^^ that's correct starts with the power rule \[\huge\rm y= x^n \rightarrow y=nx^{n-1}\] and since it's a composition function you have to take the derivative of the inner function \[y=(\color{ReD}{(x+5)^5 -1})^4\] \[y=(\color{ReD}{(\color{blue}{x+5})^5 -1})^4\]

OpenStudy (3mar):

\[y=u^4\] and \[u=g^5-1\] and \[g=x+5\] I think it would be simple for you rather than differentiating it as one shot!

OpenStudy (juan1857):

so -1 will become 0?

Nnesha (nnesha):

yes the derivative of constant is 0

OpenStudy (3mar):

\[\frac{ dy }{ dx }=\frac{ dy }{ du }*\frac{ du }{ dg }*\frac{ dg }{ dx }\] chain rule as all know

OpenStudy (juan1857):

so it would be \[y=20(x+5)^{4}(5)\]

OpenStudy (juan1857):

100(x+5)^4 ??

OpenStudy (juan1857):

wait im off

OpenStudy (juan1857):

really off

OpenStudy (juan1857):

it'' just be 20(x+5)^4 or no?

OpenStudy (mathmale):

Looking at your specific problem:\[4(((x+5)^{5})-1)^{3}\] Do the following in order: 1. apply power rule to \[4(((x+5)^{5})-1)^{3}\] Show your work, please. Then we'll discuss the next step.

OpenStudy (mathmale):

3mar: Your question is a good one, but is totally irrelevant to juan's question. Wait.

OpenStudy (juan1857):

oh i see I keep forgetting to subctract 1 from the exponent

OpenStudy (mathmale):

Apply the power rule to the given expression, please. Want to see YOUR work.

OpenStudy (juan1857):

ok let me just try to do it on paper, I was doing it in my head lol

Nnesha (nnesha):

you should keep the derivative of the outer function lets start with the easy one \[y=(\color{Red}{x^2+3})^3\] \[\rm y'=3(\color{Red}{x^2+3})^2 \cdot D~of~(\color{Red}{x^2+3})\] \[y'=3(\color{Red}{x^2+3})^2 \cdot \color{Red}{2x}\]

OpenStudy (juan1857):

I just noticed multiple errors of mine now, lol

Nnesha (nnesha):

it's impossible to do these type of questions without paper/pencil

OpenStudy (juan1857):

yeah I learned the hard way

OpenStudy (mathmale):

would you please do the entire problem, or as much of it as possible, and then share your end results. Please skip the asides (e. g., yeah, I learned the hard way).

OpenStudy (juan1857):

smh, someone closed my question and they expect me to learn

OpenStudy (3mar):

This medal for the colour, Nnesha! Thank you.

OpenStudy (juan1857):

There can someone just give me the answer at least, so that I can find my way through

OpenStudy (juan1857):

I give up on this website, its too time consuming anyways laters #Nochill @mathmale

OpenStudy (mathmale):

Sorry you're fed up. However, you must be honest with yourself: Did you actually use any of the suggestions I gave you? Did your several irrelevant remarks help you solve this problem or help anyone else to guide you? I challenge you to post a simpler function requiring application of the power and chain rules and work through this problem, and then to attempt again the one at hand.

OpenStudy (juan1857):

Truthfully I was doing your "suggestions", you don't understand since you are smart, and on my previous questions I didn't understand a word you said. How about you listen to one of my suggestions, put yourself in the shoes of another and please be more comprehensive, after all we are all human

OpenStudy (juan1857):

For the HUMBLE shall be Great, and the Great shall be Humbled